# Dolphin: Long Context as a New Modality for Energy-Efficient On-Device Language Models

### v1

@@ Abstract
This paper presents Dolphin, a novel decoder-decoder architecture for energy-efficient processing of long contexts in language models. Our approach addresses the significant energy consumption and latency challenges inherent in on-device models. Dolphin employs a compact 0.5B parameter decoder to distill extensive contextual information into a memory embedding, substantially reducing the input length for the primary 7B parameter decoder model. Inspired by vision-language models, we repurpose the image embedding projector to encode long textual contexts, effectively treating extended context as a distinct modality. This innovative method enables processing of substantially longer contexts without the typical computational overhead associated with extended input sequences. Empirical evaluations demonstrate a 10-fold improvement in energy efficiency and a 5-fold reduction in latency compared to conventional full-length context processing methods without losing quality of the response. Our work contributes to the development of more sustainable and scalable language models for on-device applications, addressing the critical need for energy-efficient and responsive AI technologies in resource-constrained environments while maintaining the accuracy to understand long contexts. This research has implications for the broader field of natural language processing, particularly in the domain of efficient model design for resource-limited settings. By enabling more sophisticated AI capabilities on edge devices, Dolphin paves the way for advanced language processing in a wide range of applications where computational resources are at a premium. The Dolphin model is publicly available at https://huggingface.co/NexaAIDev/Dolphin.
@@ 1 Introduction
In response to these challenges, we introduce Dolphin, a novel decoder-decoder architecture designed specifically for energy-efficient processing of long contexts in language models. We were inspired by recent works on Vision-Language Models (VLMs) [11, 16, 17, 18, 19], which demonstrate that model performance can benefit from specially designed multi-stage training procedures. Our approach utilizes a small 0.5B decoder to distill extensive contextual information into several memory tokens, significantly reducing the input length for the primary 7B decoder model. Drawing inspiration from vision-language models, we repurpose the image embedding projector to encode long textual contexts, effectively treating extended context as a distinct modality. This innovative method enables the processing of substantially longer contexts without incurring the typical computational overhead associated with extended input sequences. By doing so, Dolphin achieves an impressive 10-fold improvement in energy efficiency and a 5-fold improvement in latency compared to conventional full-length context processing methods. Our work contributes to the development of more sustainable and scalable language models for on-device applications, addressing the critical need for energy-efficient and fast AI technologies in resource-constrained environments while maintaining the accuracy to understand long contexts. This breakthrough has far-reaching implications for the deployment of sophisticated AI capabilities on edge devices, potentially revolutionizing fields such as mobile computing, IoT, and wearable technology.
@@ 3 Methodology
This section outlines our approach to developing the Dolphin model for efficient long-context processing. We describe the novel decoder-decoder architecture, the implementation of memory tokens, our multi-stage training process, and the dataset used for training and evaluation. These components collectively address the challenges of energy efficiency and latency in on-device language models while maintaining long-context understanding capabilities.
@@ 3.1 Long Context as a Novel Modality
In our model architecture design, we introduce an innovative decoder-decoder framework for the Dolphin model, conceptualizing long context as a novel modality. This architecture comprises two decoders of disparate sizes: a smaller decoder $\pi_{s}$ with 0.5B parameters, and a larger decoder $\pi_{l}$ with 7B parameters. The smaller decoder $\pi_{s}$ serves to transform information from the extensive context, while the larger decoder $\pi_{l}$ primarily focuses on comprehending and generating responses to the current query. Figure 1 illustrates this architecture.
Figure 1: The model architecture of the Dolphin model, with three different components, the text encoder which is a model with transformer decoder architecture. The projector is to convert the embedding information after text encoder into the embedding that can be understood by the main LLM, which is another transformers decoder model.
@@ 3.3 Multi-stage Training
Our training process for the Dolphin model comprises three distinct stages: restoration training, continual training, and instruction fine-tuning. This multi-stage approach is designed to progressively enhance the model’s ability to handle long contexts and generate appropriate responses.
@@ 3.3.4 Comparison with Vision-Language Model Training
Training Stage Dolphin (Our Model) LLaVA[11]
@@
Table 1: Comparison of training stages between Dolphin and LLaVA
@@ 3.4 Dataset
For the training and evaluation of our Dolphin model, we curated a diverse and comprehensive dataset tailored to each stage of our multi-stage training process. The dataset was designed to enhance the model’s capacity to handle long contexts, generate coherent continuations, and respond accurately to user queries across various domains. For the restoration training stage, we compiled a dataset of 100K context samples sourced from diverse domains to ensure broad coverage and generalizability. The continual training stage utilized an additional 100K context samples, distinct from those used in the restoration training, specifically curated to facilitate the model’s learning of coherent context continuation.
@@ 4.1 Testing Datasets
The testing dataset comprises 3,740 (context, prompt, response) samples derived from the Prompt-with-Context (PWC) dataset, introduced in the ICAE paper [14]. The original PWC dataset contains 240,000 samples for training and 18,000 samples for testing. We extracted 3,740 samples from the test set, selecting those with context lengths less than 512 words to align with the default maximum context length of the Dolphin model.
@@
Table 2 provides examples for each category of questions. Our testing experiments covers all six categories of questions in order to provide a comprehensive evaluation of the Dolphin model.
@@ 4.2 Restoration performance
We first evaluate the autoencoding performance of the Dolphin model during the initial training phase. This evaluation focuses on the model’s ability to accurately restore text after compression, which is crucial for maintaining semantic integrity in downstream tasks. Table 3 presents a specific example of the ICAE performing text restoration. The example demonstrates the model’s high restoration accuracy, with only a single word difference between the original and restored texts. Notably, the rare word "stuttered" is restored as "stalled," a semantically similar term that maintains the overall meaning of the passage.
@@ 4.3 Compression Performance
In this section, we present a comprehensive evaluation of the Dolphin model’s latency and compression quality compared to other candidate models. All experiments were conducted using a single NVIDIA A100 80GB GPU on the Microsoft Azure Cloud platform.
To assess the latency of the Dolphin model, which incorporates two decoders—Qwen2-0.5B and Qwen2-7B, we employed the Qwen2-7B model as a baseline. The Dolphin model leverages Qwen2-0.5B to generate compression tokens, which are then passed along with the tokenized question prompt to Qwen2-7B. In contrast, the Qwen2-7B model processes the raw input directly without any form of compression.
Our experimental results reveal that the Dolphin model significantly outperforms the Qwen2-7B model in terms of latency. As shown in Table 5, the Dolphin model achieves an average latency of 4.32 seconds, which is approximately 4.79 times faster than the Qwen2-7B model’s latency of 20.71 seconds. This substantial reduction in latency highlights the efficiency of our decoder-decoder architecture, particularly in handling long contexts with minimized computational overhead. The inclusion of a smaller 0.5B decoder for context compression is pivotal in reducing the input length for the primary 7B decoder, thereby optimizing the overall inference process.
@@
Average inference time (s) by Dolphin 4.32s
@@
For the compression quality benchmark, we utilized GPT-4[40] to evaluate the accuracy of the model responses, given the input prompt and question. The correctness scores across various question categories, as detailed in Table 4, further underscore the robustness of the Dolphin model in maintaining high accuracy while enhancing efficiency. Notably, the model achieves perfect accuracy rates of 100% in the ’Title / Keywords’ and ’Continuation’ categories, with high scores in other categories, such as 97.76% for ’Contextual QA’ and 98.53% for ’Numeric QA’. It is particularly noteworthy that even for numeric questions, which have stringent accuracy requirements due to the necessity of correct values, our model achieves an accuracy rate exceeding 98%. These results demonstrate that the Dolphin model not only excels in reducing latency but also preserves the semantic integrity and correctness of the generated outputs across a wide range of tasks.
We compare the Dolphin model with AutoCompressor [12], based on Llama-2-7b, and Qwen2-7B, the base model for Dolphin’s decoder component (Table 6). Our evaluation suggests that AutoCompressor may overfit to its training datasets, while Dolphin shows consistent performance. Notably, despite using compressed tokens, Dolphin demonstrates comparable performance to Qwen2-7B, winning 23.6% of comparisons and tying 44.2%, for a combined win-tie rate of 67.8%. This parity is significant, as compression models typically show performance degradation. The results indicate that Dolphin’s compression techniques effectively preserve, and potentially enhance, the capabilities of Qwen2-7B, while reducing computational requirements. This achievement underscores the efficacy of our approach in maintaining model performance despite information loss from token compression.
@@
The Dolphin model’s superior performance in both latency and accuracy underscores its potential for energy-efficient, on-device language modeling, especially in resource-constrained environments where balancing speed and accuracy is crucial. These findings suggest that our proposed architecture offers a promising solution for applications that demand rapid and accurate natural language processing capabilities.
@@ 5 Conclusion
In this paper, we introduced Dolphin, a novel decoder-decoder architecture designed for efficient processing of long contexts in on-device language models. By treating extended context as a distinct modality, Dolphin utilizes a compact 0.5B parameter decoder to distill contextual information into memory tokens, which are then processed by a larger 7B parameter decoder. Our experiments demonstrate that this approach achieves a 10-fold improvement in energy efficiency and a 5-fold reduction in latency compared to conventional methods, while maintaining high accuracy across various task categories.
Dolphin represents a significant advancement towards more sustainable and scalable language models for resource-constrained environments. Its multi-stage training process, comprising restoration training, continual training, and instruction fine-tuning, enables effective handling of diverse long-context tasks. Future work could explore further optimizations and adaptations of this architecture to other modalities or specialized domains.

### v2

# Squid: Long Context as a New Modality for Energy-Efficient On-Device Language Models

@@ Abstract
This paper presents Squid, a novel decoder-decoder architecture for energy-efficient processing of long contexts in language models. Our approach addresses the significant energy consumption and latency challenges inherent in on-device models. Squid employs a compact 0.5B parameter decoder to distill extensive contextual information into a memory embedding, substantially reducing the input length for the primary 7B parameter decoder model. Inspired by vision-language models, we repurpose the image embedding projector to encode long textual contexts, effectively treating extended context as a distinct modality. This innovative method enables processing of substantially longer contexts without the typical computational overhead associated with extended input sequences. Empirical evaluations demonstrate a 10-fold improvement in energy efficiency and a 5-fold reduction in latency compared to conventional full-length context processing methods without losing quality of the response. Our work contributes to the development of more sustainable and scalable language models for on-device applications, addressing the critical need for energy-efficient and responsive AI technologies in resource-constrained environments while maintaining the accuracy to understand long contexts. This research has implications for the broader field of natural language processing, particularly in the domain of efficient model design for resource-limited settings. By enabling more sophisticated AI capabilities on edge devices, Squid paves the way for advanced language processing in a wide range of applications where computational resources are at a premium. The Squid model is publicly available at https://huggingface.co/NexaAIDev/Squid.
@@ 1 Introduction
In response to these challenges, we introduce Squid, a novel decoder-decoder architecture designed specifically for energy-efficient processing of long contexts in language models. We were inspired by recent works on Vision-Language Models (VLMs) [11, 16, 17, 18, 19], which demonstrate that model performance can benefit from specially designed multi-stage training procedures. Our approach utilizes a small 0.5B decoder to distill extensive contextual information into several memory tokens, significantly reducing the input length for the primary 7B decoder model. Drawing inspiration from vision-language models, we repurpose the image embedding projector to encode long textual contexts, effectively treating extended context as a distinct modality. This innovative method enables the processing of substantially longer contexts without incurring the typical computational overhead associated with extended input sequences. By doing so, Squid achieves an impressive 10-fold improvement in energy efficiency and a 5-fold improvement in latency compared to conventional full-length context processing methods. Our work contributes to the development of more sustainable and scalable language models for on-device applications, addressing the critical need for energy-efficient and fast AI technologies in resource-constrained environments while maintaining the accuracy to understand long contexts. This breakthrough has far-reaching implications for the deployment of sophisticated AI capabilities on edge devices, potentially revolutionizing fields such as mobile computing, IoT, and wearable technology.
@@ 3 Methodology
This section outlines our approach to developing the Squid model for efficient long-context processing. We describe the novel decoder-decoder architecture, the implementation of memory tokens, our multi-stage training process, and the dataset used for training and evaluation. These components collectively address the challenges of energy efficiency and latency in on-device language models while maintaining long-context understanding capabilities.
@@ 3.1 Long Context as a Novel Modality
In our model architecture design, we introduce an innovative decoder-decoder framework for the Squid model, conceptualizing long context as a novel modality. This architecture comprises two decoders of disparate sizes: a smaller decoder $\pi_{s}$ with 0.5B parameters, and a larger decoder $\pi_{l}$ with 7B parameters. The smaller decoder $\pi_{s}$ serves to transform information from the extensive context, while the larger decoder $\pi_{l}$ primarily focuses on comprehending and generating responses to the current query. Figure 1 illustrates this architecture.
Figure 1: The model architecture of the Squid model, with three different components, the text encoder which is a model with transformer decoder architecture. The projector is to convert the embedding information after text encoder into the embedding that can be understood by the main LLM, which is another transformers decoder model.
@@ 3.3 Multi-stage Training
Our training process for the Squid model comprises three distinct stages: restoration training, continual training, and instruction fine-tuning. This multi-stage approach is designed to progressively enhance the model’s ability to handle long contexts and generate appropriate responses.
@@ 3.3.4 Comparison with Vision-Language Model Training
Training Stage Squid (Our Model) LLaVA[11]
@@
Table 1: Comparison of training stages between Squid and LLaVA
@@ 3.4 Dataset
For the training and evaluation of our Squid model, we curated a diverse and comprehensive dataset tailored to each stage of our multi-stage training process. The dataset was designed to enhance the model’s capacity to handle long contexts, generate coherent continuations, and respond accurately to user queries across various domains. For the restoration training stage, we compiled a dataset of 100K context samples sourced from diverse domains to ensure broad coverage and generalizability. The continual training stage utilized an additional 100K context samples, distinct from those used in the restoration training, specifically curated to facilitate the model’s learning of coherent context continuation.
@@ 4.1 Testing Datasets
The testing dataset comprises 3,740 (context, prompt, response) samples derived from the Prompt-with-Context (PWC) dataset, introduced in the ICAE paper [14]. The original PWC dataset contains 240,000 samples for training and 18,000 samples for testing. We extracted 3,740 samples from the test set, selecting those with context lengths less than 512 words to align with the default maximum context length of the Squid model.
@@
Table 2 provides examples for each category of questions. Our testing experiments covers all six categories of questions in order to provide a comprehensive evaluation of the Squid model.
@@ 4.2 Restoration performance
We first evaluate the autoencoding performance of the Squid model during the initial training phase. This evaluation focuses on the model’s ability to accurately restore text after compression, which is crucial for maintaining semantic integrity in downstream tasks. Table 3 presents a specific example of the ICAE performing text restoration. The example demonstrates the model’s high restoration accuracy, with only a single word difference between the original and restored texts. Notably, the rare word "stuttered" is restored as "stalled," a semantically similar term that maintains the overall meaning of the passage.
@@ 4.3 Compression Performance
In this section, we present a comprehensive evaluation of the Squid model’s latency and compression quality compared to other candidate models. All experiments were conducted using a single NVIDIA A100 80GB GPU on the Microsoft Azure Cloud platform.
To assess the latency of the Squid model, which incorporates two decoders—Qwen2-0.5B and Qwen2-7B, we employed the Qwen2-7B model as a baseline. The Squid model leverages Qwen2-0.5B to generate compression tokens, which are then passed along with the tokenized question prompt to Qwen2-7B. In contrast, the Qwen2-7B model processes the raw input directly without any form of compression.
Our experimental results reveal that the Squid model significantly outperforms the Qwen2-7B model in terms of latency. As shown in Table 5, the Squid model achieves an average latency of 4.32 seconds, which is approximately 4.79 times faster than the Qwen2-7B model’s latency of 20.71 seconds. This substantial reduction in latency highlights the efficiency of our decoder-decoder architecture, particularly in handling long contexts with minimized computational overhead. The inclusion of a smaller 0.5B decoder for context compression is pivotal in reducing the input length for the primary 7B decoder, thereby optimizing the overall inference process.
@@
Average inference time (s) by Squid 4.32s
@@
For the compression quality benchmark, we utilized GPT-4[40] to evaluate the accuracy of the model responses, given the input prompt and question. The correctness scores across various question categories, as detailed in Table 4, further underscore the robustness of the Squid model in maintaining high accuracy while enhancing efficiency. Notably, the model achieves perfect accuracy rates of 100% in the ’Title / Keywords’ and ’Continuation’ categories, with high scores in other categories, such as 97.76% for ’Contextual QA’ and 98.53% for ’Numeric QA’. It is particularly noteworthy that even for numeric questions, which have stringent accuracy requirements due to the necessity of correct values, our model achieves an accuracy rate exceeding 98%. These results demonstrate that the Squid model not only excels in reducing latency but also preserves the semantic integrity and correctness of the generated outputs across a wide range of tasks.
We compare the Squid model with AutoCompressor [12], based on Llama-2-7b, and Qwen2-7B, the base model for Squid’s decoder component (Table 6). Our evaluation suggests that AutoCompressor may overfit to its training datasets, while Squid shows consistent performance. Notably, despite using compressed tokens, Squid demonstrates comparable performance to Qwen2-7B, winning 23.6% of comparisons and tying 44.2%, for a combined win-tie rate of 67.8%. This parity is significant, as compression models typically show performance degradation. The results indicate that Squid’s compression techniques effectively preserve, and potentially enhance, the capabilities of Qwen2-7B, while reducing computational requirements. This achievement underscores the efficacy of our approach in maintaining model performance despite information loss from token compression.
@@
The Squid model’s superior performance in both latency and accuracy underscores its potential for energy-efficient, on-device language modeling, especially in resource-constrained environments where balancing speed and accuracy is crucial. These findings suggest that our proposed architecture offers a promising solution for applications that demand rapid and accurate natural language processing capabilities.
@@ 5 Conclusion
In this paper, we introduced Squid, a novel decoder-decoder architecture designed for efficient processing of long contexts in on-device language models. By treating extended context as a distinct modality, Squid utilizes a compact 0.5B parameter decoder to distill contextual information into memory tokens, which are then processed by a larger 7B parameter decoder. Our experiments demonstrate that this approach achieves a 10-fold improvement in energy efficiency and a 5-fold reduction in latency compared to conventional methods, while maintaining high accuracy across various task categories.
Squid represents a significant advancement towards more sustainable and scalable language models for resource-constrained environments. Its multi-stage training process, comprising restoration training, continual training, and instruction fine-tuning, enables effective handling of diverse long-context tasks. Future work could explore further optimizations and adaptations of this architecture to other modalities or specialized domains.